OpenStudy (anonymous):

If A is IaIa, Iai and B is is IbIb, Ibi, What is the phenotype and the genotype of a couple who's blood type is hetero A and Hetero B?

OpenStudy (abhisar):

\(\sf \huge A=I^aI^a~and~I^ai\\\huge B=I^bI^b~and~I^bi\) \(\sf Hetero~A=I^ai\\Hetero~B=I^bi\) Following will be the cross between them